A psychologist is studying the capacity of the human mind to process information. He has found that the percentage of information one particular participant can recall after t months have passed can be modeled by the following logarithmic function.

f(t)=92−22ln(t+1), where 1≤t≤12
What percentage of information is retained after 8 months? Round your answer to the nearest whole percentage, if necessary.
1
Expert's answer
2020-11-16T19:49:17-0500

The percentage of information that is retained after "8" months is


"f(8)=92-22\ln(8+1)=92-22\ln 9=92-22\ln 3^2=92-44\ln 3=43.661."


Answer: "44"